# Diffusion Models for Image Restoration and Enhancement: A Comprehensive Survey

### v2

@@ 3.3.4 Generate Residuals
In the supervised DM-based IR, most studies directly generate high-quality images from noise, which necessitates the concurrent generation of structure and textures. However, regenerating the structures/textures already existing in low-quality images unnecessarily burdens the diffusion model and increases extra resource costs. Motivated by this, some representative studies (Li et al., 2022b; Whang et al., 2022; Shang et al., 2023a; Liu et al., 2024b; Yue et al., 2023; Gao and Shan, 2022; Karchev et al., 2022; Liu et al., 2023a) seek to move the generation process from the image space to the residual space. The objective is to generate the residuals between paired high-quality and low-quality images. As the pioneering work, SRDiff (Li et al., 2022b) is the first to utilize the diffusion model to predict the residual in SR. Whang (Whang et al., 2022) introduces a predict-and-refine strategy for image deblurring tasks. As shown in Fig. 7, this work first predicts an initial deblurring image with the deterministic deblurring network and then generates the residual through a stochastic diffusion model. ResDiff (Shang et al., 2023a) mentioned above also adopts this strategy for SR. Resshift (Yue et al., 2023) introduces a novel Markov chain that shifts residuals between high-quality and low-quality images. This residual shifting technique achieves comparable performance even with a reduced number of sampling steps, as few as 15. Another work RDDM (Liu et al., 2024b) redefines the forward diffusion process and incorporates residuals to represent directional diffusion from the target domain to the input domain. RDDM predicts both noise and residuals during training to effectively restore images.
@@ 3.3.5 Diffusion on Latent Space
To alleviate the training and sampling costs of the diffusion model, some works (Rombach et al., 2022; Luo et al., 2023c; Xia et al., 2023; Yang et al., 2024a; Chung et al., 2023c; Qu et al., 2024; Luo et al., 2023a; Sun et al., 2023), try to conduct diffuion in latent space. Among them, stableDiffusion (Rombach et al., 2022) is the first work to implement the DM-based generation in latent space. Particularly, it pre-trains an autoencoding model (i.e., an encoder-decoder architecture) to learn the perceptual latent space, which is able to preserve the perceptual quality of the reconstructed image while reducing the computational complexity. Utilizing the pre-trained autoencoder, StableDiffusion transforms the image-wise diffusion process to the latent space, and then introduces various conditions (e.g., text, segmentation map, and image) into the diffusion model with the cross-attention mechanism. Inspired by this, Refusion (Luo et al., 2023c) introduces the latent-wise diffusion model for image restoration to accelerate the training and sampling, which is shown in Fig. 8. In contrast to the above works, where latent space is obtained by compressing the original image, DiffIR (Xia et al., 2023) exploits the latent-wise diffusion model to generate the compact IR priors, which guides dynamic transformer-based restoration network (DIRformer) to achieve better restoration.
Figure 8: An overview of diffusion model used in Refusion (Luo et al., 2023c), the low-quality image is first compressed into latent by an encoder. Then the latent information is modeled by a diffusion process. Here, $F$ denotes the forward process, which progressively adds noise to $x_{0}$, transforming it into $x_{T}$.
@@ 3.3.6 Efficient Transfer-Learning
In supervised DM-based IR, efficient and rapid fine-tuning has become a focal point of current research (Wang et al., 2023c; Lin et al., 2024; Yang et al., 2024a; Yu et al., 2024; Jiang et al., 2024b; Ai et al., 2024). This method principally leverages pre-trained models such as DDPM or Stable Diffusion, extracting certain distortion information through feature extraction and subsequently modulating it into the diffusion model. Pioneering work is StableSR, where a time-aware encoder is trained to extract information from low-quality images and incorporate the learned features into pretrained stable diffusion(trained on LAION-5B (Schuhmann et al., 2022)). On the other hand, DiffBIR (Lin et al., 2024) and PASD (Yang et al., 2024b) introduce conditions by training a UNet encoder using zero convolutions, which bears similarities to ControlNet (Zhang and Agrawala, 2023). Another significant advancement in this area is SUPIR (Yu et al., 2024), which builds upon a pre-trained model based on SDXL (Podell et al., 2023). It extracts textual prompts through LLaVA (Liu et al., 2024a) and introduces distorted image information using a trimmed Controlnet.
@@
To mitigate the primary challenges in zero-shot DM-based IR, the projection-based method has been introduced in some studies (Lugmayr et al., 2022; Choi et al., 2021; Chung et al., 2022b; Moser et al., 2023b). This approach aims to extract inherent structures/textures from low-quality images as complementary to generated images at each step, which can ensure data consistency. For instance, the task of image inpainting involves merely generating the content for the mask region. The unmasked region of the low-quality image can substitute the corresponding part of the denoised image at the t1t-1 step, thereby establishing the condition for data consistency during the sampling process. In line with this, RePaint (Lugmayr et al., 2022) exploits a simple projection for image inpainting task:
@@
where $x_{t-1}^{known}\sim\mathcal{N}(\sqrt{\hat{\alpha}_{t}}y,1-\hat{\alpha}_{t}\mathbf{I})$ is the diffused results at time step t1t-1 by adding noise for masked image $y$. $x_{t-1}^{unknown}$ is sampled from denoised prediction of diffusion model. In contrast, ILVR (Choi et al., 2021) employs the low-frequency projection for image super-resolution. Theoretically, at time step t1t-1, the predicted latent variable $x_{t-1}$ and $y_{t-1}$ (i.e., adding noise to low-resolution image $y$ at the t1t-1 step in the diffusion process) should share the same low-frequency components. Consequently, it substitutes the low-frequency components of $x_{t-1}$ with its counterpart from $y_{t-1}$, ensuring data consistency and establishing an improved condition for the diffusion model. As an advanced solution, CCDF (Chung et al., 2022b) introduces a unified projection method as:
@@
where $H$ is the linear degradation operator and $z$ is a contaminating noise. In this setting, the condition probability $p(x|y)$ cannot be directly estimated since the existence of noise $z$. To get rid of the noise $z$, SNIPS (Kawar et al., 2021) and DDRM (Kawar et al., 2022a) run the diffusion process in the spectral domain with singular value decomposition (SVD) on the degradation operator $H$. In particular, SNIPS (Kawar et al., 2021) is based on the annealed Langevian dynamics and derives the conditional score function on the spectral space, which achieves great performance on image deblurring, super-resolution, and compressive sensing tasks. Following SNIPS, DDRM (Kawar et al., 2022a) further extends the SVD decomposition to the variational objective of linear reverse problems, which reveals a pre-trained DDPM (Ho et al., 2020)/DDIM (Song et al., 2020a) can be the optimal solution for it. Notably, the above works only focus on the linear reverse problem. In contrast, Kawar et al (Kawar et al., 2022b) investigate non-linear inverse problems based on the special case of DDRM (i.e., no noise $z$ in the reverse problem), and extend the pseudo-inverse concept to achieve JPEG artifact correction. For the MRI reconstruction, SVD decomposition is not suitable. To overcome this, Song et al. train the unconditional generative model for medical images from scratch and then exploit the matrix decomposition in the sampling process to solve the linear reverse problem, which is general for unknown measurement processes.
With a different purpose, another decomposition strategy, range-null space decomposition is introduced by DDNM (Wang et al., 2023g) to further improve the zero-shot image restoration, where the range space is responsible for the data consistency, and null space is used to improve the reality (i.e., the perceptual quality). Given the noiseless inverse $y=Hx$, it can be decomposed into:
@@
where $H^{\dagger}$ is the pseudo-inverse of degradation operation $H$. We can see the range space $HH^{\dagger}Hx=Hx=y$ can ensure data consistency, and null space $H(I-H^{\dagger}H)x$ has no effects on data consistency since $H(I-H^{\dagger}H)x=\mathbf{0}$. As shown in Fig. 9, based on this, DDNM (Wang et al., 2023g) rectifies the prediction of $x_{0}$ at time step $t$ as: $\hat{x}_{0|t}=H^{\dagger}y+(I-H^{\dagger}H)x_{0|t}$, where $x_{0|t}$ can be estimated with the noise prediction at time step $t$. With the rectified $\hat{x}_{0|t}$, we can compute the denoised output $x_{t-1}$ at time step t1t-1, which ensures the data consistency and serves as a better condition for the next noise prediction. Moreover, DDNM also exploits the SVD to solve the linear inverse problem with noise, termed DDNM+.
Figure 9: The architecture of DDNM (Wang et al., 2023g),the forward degradation operator AA is decomposed into range-null space to meet the realness and data consistency.
@@
The projection-based methods have shown superior performance in the inverse problem of image restoration, where a projection-based measurement consistency correction is added after the reverse sampling step of the diffusion model. However, most projection-based works are devoted to the noiseless inverse problems and usually suffer from unsatisfied data consistency, since the projection throws the sample path off the data manifold (Chung et al., 2022a). To solve the general noisy linear reverse problem, some works (Chung et al., 2022a, 2023b; Fabian et al., 2023; Song et al., 2023; Zhang et al., 2023b; Fei et al., 2023; Cardoso et al., 2023; Garber and Tirer, 2024; Lee et al., 2024; Cao et al., 2024) aim to estimate the posterior distribution $p(x|y)$ with the unconditional diffusion model based on the Bayes theorem. It is equivalent to estimating the conditional posterior $p(x_{t}|y)$ at each step of the reverse process. Based on the Bayes theorem, it can be derived as:
@@
As the pioneering works, MCG (Chung et al., 2022a) and DPS (Chung et al., 2023b) approximate the posterior $p(y|x_{t})$ with $p(y|\hat{x}_{o})$, and $\hat{x}_{o}$ is the expectation given $x_{t}$ as $\hat{x}_{o}=E[x_{o}|x_{t}]$ (Chung et al., 2023b) with Tweedie’s formula. Concretely, MCG (Chung et al., 2022a) considers the data consistency from the perspective of the data manifold, where the manifold constrained gradient is proposed to let the correction lie on the data manifold. However, DPS (Chung et al., 2023b) pinpoints that the projection operation in MCG is harmful to data consistency since it might cause the sampling path off the data manifold. Based on this, DPS (Chung et al., 2023b) discards the projection step in the reverse process and estimates the posterior as:
@@
Different from the above works, some works (Fei et al., 2023; Zhang et al., 2023b) attempt to model $p(y|x_{t})$ with other strategies. It is noteworthy that the higher conditional probability $p(y|x_{t})$ is equivalent to a smaller distance between $D(x_{t})$ and $y$ (Fei et al., 2023). Therefore, GDP (Fei et al., 2023) proposes a heuristic approximation of distribution $p(y|x_{t})$ as follows:
@@
where $\mathcal{L}$ and $\mathcal{Q}$ denote the distance metric and the quality loss, respectively. $Z$ is a normalization factor and $s$ is the scaling factor that controls the weight of guidance. However, the distance $\mathcal{L}$ is hard to be defined since the noise magnitudes in $x_{t}$ and $y$ are different. Thus, they substitute $x_{t}$ with its clean estimation $\hat{x}_{0}$ in the distance measurement. With the same purpose, Copaint (Zhang et al., 2023b) tries to predict the $\hat{x}_{0}$ through the one-step estimation by a neural network.
Instead of modeling intractable distribution $p(y|x_{t})$, Feng et al.(Feng et al., 2023) directly estimates the posterior $p(x_{t}|y)$ from a variational perspective. Following DPI (Sun and Bouman, 2021; Sun et al., 2022), they define a family of distributions $q_{\theta}$ through RealNVP (Dinh et al., 2016) normalizing flow with parameter $\theta$, which is optimized through a minimal KL-divergence between true posterior and estimated distribution $q_{\theta}$.
@@ 4 Diffusion Models for Blind/Real-world Image Restoration
Although the methods in Sec. 3 have achieved great breakthroughs in image restoration, most of them (Chung et al., 2023b; Saharia et al., 2022b; Kawar et al., 2022a; Li et al., 2022b; Saharia et al., 2022a; Özdenizci and Legenstein, 2023; Ho et al., 2022; Song et al., 2023; Wang et al., 2023g) focus on solving the synthetic distortions, which usually perform poorly in the out-of-distribution (OOD) real-world/blind degradations. The reasons stem from the inherent challenges of real-world IR: 1) the unknown degradation modes are hard to be identified. 2) collecting distorted/clean image pairs is non-trivial and even unavailable in the real world. To overcome this, previous works (Wang et al., 2022c; Liang et al., 2022; Wang et al., 2022d) have attempted to solve it by simulating real-world degradations (Wang et al., 2021b; Zhang et al., 2021b, 2022b; Wang et al., 2022c), and unsupervised learning (Wang et al., 2021a; Lugmayr et al., 2019a; Wang et al., 2022d), etc. Inspired by these, some pioneering works (Sahak et al., 2023; Chung et al., 2023a; Wang et al., 2023i; Yin et al., 2022; Abu-Hussein et al., 2022) begin to explore how to exploit diffusion models to solve real-world degradations. In this paper, we divided the DM-based blind/real-world IR (Sahak et al., 2023; Yin et al., 2022; Ren et al., 2022; Wang et al., 2023c; Welker et al., 2022; Yang et al., 2023b; Chan et al., 2023; Wang et al., 2023i; Murata et al., 2023; Chung et al., 2023a; Miao et al., 2023) into four categories, i.e., distortion simulation (Sahak et al., 2023; Yang et al., 2023b), kernel estimation (Chung et al., 2023a; Murata et al., 2023), domain translation (Wei et al., 2023; Yang et al., 2023b), and Distortion-invariant diffusion model (Wang et al., 2023i; Yue and Loy, 2024; Ren et al., 2022).

### v3

@@ 3.3.4 Generate Residuals
In the supervised DM-based IR, most studies directly generate high-quality images from noise, which necessitates the concurrent generation of structure and textures. However, regenerating the structures/textures already existing in low-quality images unnecessarily burdens the diffusion model and increases extra resource costs. Motivated by this, some representative studies (li2022srdiff, ; deblur-DPM, ; shang2023resdiff, ; liu2023residual-RDDM, ; yue2023resshift, ; gao2022cocodiff-residual2, ; karchev2022strong-residual1, ; liu2023textdiff, ) seek to move the generation process from the image space to the residual space. The objective is to generate the residuals between paired high-quality and low-quality images. As the pioneering work, SRDiff (li2022srdiff, ) is the first to utilize the diffusion model to predict the residual in SR. Whang (deblur-DPM, ) introduces a predict-and-refine strategy for image deblurring tasks. As shown in Fig. 7, this work first predicts an initial deblurring image with the deterministic deblurring network and then generates the residual through a stochastic diffusion model. ResDiff (shang2023resdiff, ) mentioned above also adopts this strategy for SR. Resshift (yue2023resshift, ) introduces a novel Markov chain that shifts residuals between high-quality and low-quality images. This residual shifting technique achieves comparable performance even with a reduced number of sampling steps, as few as 15. Another work RDDM (liu2023residual-RDDM, ) redefines the forward diffusion process and incorporates residuals to represent directional diffusion from the target domain to the input domain. RDDM predicts both noise and residuals during training to effectively restore images.
@@ 3.3.5 Diffusion on Latent Space
To alleviate the training and sampling costs of the diffusion model, some works (LDM, ; luo2023refusion, ; xia2023diffir, ; yang2023pixel-aware, ; chung2023prompt-P2L, ; qu2024xpsr, ; luo2023image-frequency-feng, ; sun2023improving-CCSR, ), try to conduct diffuion in latent space. Among them, stableDiffusion (LDM, ) is the first work to implement the DM-based generation in latent space. Particularly, it pre-trains an autoencoding model (i.e., an encoder-decoder architecture) to learn the perceptual latent space, which is able to preserve the perceptual quality of the reconstructed image while reducing the computational complexity. Utilizing the pre-trained autoencoder, StableDiffusion transforms the image-wise diffusion process to the latent space, and then introduces various conditions (e.g., text, segmentation map, and image) into the diffusion model with the cross-attention mechanism. Inspired by this, Refusion (luo2023refusion, ) introduces the latent-wise diffusion model for image restoration to accelerate the training and sampling, which is shown in Fig. 8. In contrast to the above works, where latent space is obtained by compressing the original image, DiffIR (xia2023diffir, ) exploits the latent-wise diffusion model to generate the compact IR priors, which guides dynamic transformer-based restoration network (DIRformer) to achieve better restoration.
Figure 8: An overview of diffusion model used in Refusion (luo2023refusion, ), the low-quality image is first compressed into latent by an encoder. Then the latent information is modeled by a diffusion process. Here, $F$ denotes the forward process, which progressively adds noise to $x_{0}$, transforming it into $x_{T}$.
@@ 3.3.6 Efficient Transfer-Learning
In supervised DM-based IR, efficient and rapid fine-tuning has become a focal point of current research (wang2023exploiting-stable-sr, ; lin2023diffbir, ; yang2023pixel-aware, ; yu2024scaling-SUPIR, ; jiang2023autodir, ; ai2023multimodal-mperceriver, ). This method principally leverages pre-trained models such as DDPM or Stable Diffusion, extracting certain distortion information through feature extraction and subsequently modulating it into the diffusion model. Pioneering work is StableSR, where a time-aware encoder is trained to extract information from low-quality images and incorporate the learned features into pretrained stable diffusion(trained on LAION-5B (schuhmann2022laion-5B, )). On the other hand, DiffBIR (lin2023diffbir, ) and PASD (yang2023pixel-PASD, ) introduce conditions by training a UNet encoder using zero convolutions, which bears similarities to ControlNet (zhang2023adding-ControlNet, ). Another significant advancement in this area is SUPIR (yu2024scaling-SUPIR, ), which builds upon a pre-trained model based on SDXL (podell2023sdxl, ). It extracts textual prompts through LLaVA (liu2024visual-LLaVA, ) and introduces distorted image information using a trimmed Controlnet.
@@
To mitigate the primary challenges in zero-shot DM-based IR, the projection-based method has been introduced in some studies (repaint, ; choi2021ilvr, ; CCDF, ; moser2023yoda, ). This approach aims to extract inherent structures/textures from low-quality images as complementary to generated images at each step, which can ensure data consistency. For instance, the task of image inpainting involves merely generating the content for the mask region. The unmasked region of the low-quality image can substitute the corresponding part of the denoised image at the t1t-1 step, thereby establishing the condition for data consistency during the sampling process. In line with this, RePaint (repaint, ) exploits a simple projection for image inpainting task:
@@
where $x_{t-1}^{known}\sim\mathcal{N}(\sqrt{\hat{\alpha}_{t}}y,1-\hat{\alpha}_{t}\mathbf{I})$
is the diffused results at time step t1t-1 by adding noise for masked image $y$. $x_{t-1}^{unknown}$ is sampled from denoised prediction of diffusion model. In contrast, ILVR (choi2021ilvr, ) employs the low-frequency projection for image super-resolution.
Theoretically, at time step t1t-1, the predicted latent variable $x_{t-1}$ and $y_{t-1}$ (i.e., adding noise to low-resolution image $y$ at the t1t-1 step in the diffusion process) should share the same low-frequency components. Consequently, it substitutes the low-frequency components of $x_{t-1}$ with its counterpart from $y_{t-1}$, ensuring data consistency and establishing an improved condition for the diffusion model. As an advanced solution, CCDF (CCDF, ) introduces a unified projection method as:
@@
where $H$ is the linear degradation operator and $z$ is a contaminating noise. In this setting, the condition probability $p(x|y)$ cannot be directly estimated since the existence of noise $z$. To get rid of the noise $z$, SNIPS (kawar2021snips, ) and DDRM (DDRMDRM, ) run the diffusion process in the spectral domain with singular value decomposition (SVD) on the degradation operator $H$. In particular, SNIPS (kawar2021snips, ) is based on the annealed Langevian dynamics and derives the conditional score function on the spectral space, which achieves great performance on image deblurring, super-resolution, and compressive sensing tasks. Following SNIPS, DDRM (DDRMDRM, ) further extends the SVD decomposition to the variational objective of linear reverse problems, which reveals a pre-trained DDPM (DDPM, )/DDIM (DDIM, ) can be the optimal solution for it. Notably, the above works only focus on the linear reverse problem. In contrast, Kawar et al (jpegddrm, ) investigate non-linear inverse problems based on the special case of DDRM (i.e., no noise $z$ in the reverse problem), and extend the pseudo-inverse concept to achieve JPEG artifact correction. For the MRI reconstruction, SVD decomposition is not suitable. To overcome this, Song et al. train the unconditional generative model for medical images from scratch and then exploit the matrix decomposition in the sampling process to solve the linear reverse problem, which is general for unknown measurement processes.
With a different purpose, another decomposition strategy, range-null space decomposition is introduced by DDNM (ddnm, ) to further improve the zero-shot image restoration, where the range space is responsible for the data consistency, and null space is used to improve the reality (i.e., the perceptual quality). Given the noiseless inverse $y=Hx$, it can be decomposed into:
@@
where $H^{\dagger}$ is the pseudo-inverse of degradation operation $H$. We can see the range space $HH^{\dagger}Hx=Hx=y$ can ensure data consistency, and null space $H(I-H^{\dagger}H)x$ has no effects on data consistency since $H(I-H^{\dagger}H)x=\mathbf{0}$. As shown in Fig. 9, based on this, DDNM (ddnm, ) rectifies the prediction of $x_{0}$ at time step $t$ as: $\hat{x}_{0|t}=H^{\dagger}y+(I-H^{\dagger}H)x_{0|t}$, where $x_{0|t}$ can be estimated with the noise prediction at time step $t$. With the rectified $\hat{x}_{0|t}$, we can compute the denoised output $x_{t-1}$ at time step t1t-1, which ensures the data consistency and serves as a better condition for the next noise prediction. Moreover, DDNM also exploits the SVD to solve the linear inverse problem with noise, termed DDNM+.
Figure 9: The architecture of DDNM (ddnm, ),the forward degradation operator AA is decomposed into range-null space to meet the realness and data consistency.
@@
The projection-based methods have shown superior performance in the inverse problem of image restoration, where a projection-based measurement consistency correction is added after the reverse sampling step of the diffusion model. However, most projection-based works are devoted to the noiseless inverse problems and usually suffer from unsatisfied data consistency, since the projection throws the sample path off the data manifold (mcg, ). To solve the general noisy linear reverse problem, some works (mcg, ; dps, ; fabian2023diracdiffusion, ; iigdm, ; copaint, ; GDP, ; cardoso2023monte, ; garber2023image-DDPG, ; lee2024spatial-SAFARA, ; cao2023deep-DeqIR, ) aim to estimate the posterior distribution $p(x|y)$ with the unconditional diffusion model based on the Bayes theorem. It is equivalent to estimating the conditional posterior $p(x_{t}|y)$ at each step of the reverse process. Based on the Bayes theorem, it can be derived as:
@@
As the pioneering works, MCG (mcg, ) and DPS (dps, ) approximate the posterior $p(y|x_{t})$ with $p(y|\hat{x}_{o})$, and $\hat{x}_{o}$ is the expectation given $x_{t}$ as $\hat{x}_{o}=E[x_{o}|x_{t}]$ (dps, ) with Tweedie’s formula. Concretely, MCG (mcg, ) considers the data consistency from the perspective of the data manifold, where the manifold constrained gradient is proposed to let the correction lie on the data manifold. However, DPS (dps, ) pinpoints that the projection operation in MCG is harmful to data consistency since it might cause the sampling path off the data manifold. Based on this, DPS (dps, ) discards the projection step in the reverse process and estimates the posterior as:
@@
Different from the above works, some works (GDP, ; copaint, ) attempt to model $p(y|x_{t})$ with other strategies. It is noteworthy that the higher conditional probability $p(y|x_{t})$ is equivalent to a smaller distance between $D(x_{t})$ and $y$ (GDP, ). Therefore, GDP (GDP, ) proposes a heuristic approximation of distribution $p(y|x_{t})$ as follows:
@@
where $\mathcal{L}$ and $\mathcal{Q}$ denote the distance metric and the quality loss, respectively. $Z$ is a normalization factor and $s$ is the scaling factor that controls the weight of guidance. However, the distance $\mathcal{L}$ is hard to be defined since the noise magnitudes in $x_{t}$ and $y$ are different. Thus, they substitute $x_{t}$ with its clean estimation $\hat{x}_{0}$ in the distance measurement. With the same purpose, Copaint (copaint, ) tries to predict the $\hat{x}_{0}$ through the one-step estimation by a neural network.
Instead of modeling intractable distribution $p(y|x_{t})$, Feng et al.(feng2023score, ) directly estimates the posterior $p(x_{t}|y)$ from a variational perspective. Following DPI (sun2021deep-dpi1, ; sun2022alpha-dpi2, ), they define a family of distributions $q_{\theta}$ through RealNVP (RealNVP, ) normalizing flow with parameter $\theta$, which is optimized through a minimal KL-divergence between true posterior and estimated distribution $q_{\theta}$.
@@ 4 Diffusion Models for Blind/Real-world Image Restoration
Although the methods in Sec. 3 have achieved great breakthroughs in image restoration, most of them (dps, ; SR3, ; DDRMDRM, ; li2022srdiff, ; saharia2022palette, ; weather-diff, ; CDM, ; iigdm, ; ddnm, ) focus on solving the synthetic distortions, which usually perform poorly in the out-of-distribution (OOD) real-world/blind degradations. The reasons stem from the inherent challenges of real-world IR: 1) the unknown degradation modes are hard to be identified. 2) collecting distorted/clean image pairs is non-trivial and even unavailable in the real world. To overcome this, previous works (wang2022blind-simulate3, ; liang2022efficient-simulating2, ; wang2022ucl-unsupervised, ) have attempted to solve it by simulating real-world degradations (wang2021realesrgan, ; zhang2021designing-sr3+, ; zhang2022single-simulating1, ; wang2022blind-simulate3, ), and unsupervised learning (wang2021unsupervised, ; lugmayr2019unsupervised, ; wang2022ucl-unsupervised, ), etc. Inspired by these, some pioneering works (SR3+, ; blinddps, ; wang2023dr2, ; yin2022diffgar, ; abu2022adir, ) begin to explore how to exploit diffusion models to solve real-world degradations. In this paper, we divided the DM-based blind/real-world IR (SR3+, ; yin2022diffgar, ; dg-dpm, ; wang2023exploiting-stable-sr, ; welker2022driftrec, ; yang2023synthesizing, ; chan2023sud, ; wang2023dr2, ; murata2023gibbsddrm, ; blinddps, ; miao2023dds2m, ) into four categories, i.e., distortion simulation (SR3+, ; yang2023synthesizing, ), kernel estimation (blinddps, ; murata2023gibbsddrm, ), domain translation (wei2023raindiffusion, ; yang2023synthesizing, ), and Distortion-invariant diffusion model (wang2023dr2, ; yue2022difface, ; dg-dpm, ).